an unbanked curve has a radius of 60m. the maximum speed at which a car can make a turn if the cofficient of static friction is 0.75 is
1
Expert's answer
2012-10-04T08:27:48-0400
t=3
sec
s=60m
coff=0.75
f=coff*m*g=m*a
coff*g=a=7.5m^2/sec
s=|vt+(at^2)/2|
v=|(-s+(at^2)/2)/t|=|(-60+7.5*9/2)/3|=8.75m/sec